OpenStudy (anonymous):

sienna posted 2/3 of her collection of photographs on a page of er web site. additionally, she posted 4 photographs on the homepage of her website. the total number of photographs that she posted on her web site was 22. how many photographs does sienna have in her collection?

OpenStudy (anonymous):

So 22 is 2/3 of her collection + 4, so take away the 4 and you have 2/3 of her collection = 18!\[\frac{ 2 }{ 3 }x=18\] where x is her collection. Can you get it from there?

OpenStudy (anonymous):

no totally confused

OpenStudy (anonymous):

Haha, sorry! So, 22 is the total amount of photos they posted. Which is made up of the 2/3 of her collection, as well as the 4 on the homepage, yes?

OpenStudy (anonymous):

thank you but ahint too the answer, is it less than 22

OpenStudy (anonymous):

No, I don't think it is! Do you need help with the equation I got to before?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, \[\frac{ 2 }{ 3 }x=18\]Times both sides by 3 to get rid of the fraction\[2x=18\times3=54\]Then divide both sides by 2, so you have\[x=54/2=27\] where x is the collection!

OpenStudy (anonymous):

thank you!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

No problem!